# Clusterpath Gaussian Graphical Modeling

### v2

@@ 2.3 Cyclic Block Coordinate Descent Algorithm
We develop a cyclic block coordinate descent algorithm, tailored to minimizing objective function (2). [^1] One could opt for alternative algorithms such as the alternating directions method of multipliers (ADMM), provided it is tailored towards solving problem (2); as for instance Wilms & Bien, (2022) do for their method. However, while all the subproblems of the ADMM by Wilms & Bien, (2022) are solvable in closed-form, this would not be the case for the subproblem corresponding to our proposed aggregation penalty. This reason, in combination with the natural link between the block structure in $\bm{\Theta}$ and the block structure of the cyclic block coordinate descent algorithm further explains our choice for the latter. We opt for cyclic block coordinate descent since the blocks are naturally formed by the clusters $k=1,\ldots,K$. By exploiting this block structure in $\mathbf{\Theta}$, our algorithm builds on efficient expressions of the objective function, which allows us to minimize optimization problem (2) in a computationally efficient way.
@@
In practice, the block structure of the precision matrix is not known beforehand; our penalization problem (2) jointly identifies the block structure and estimates the precision matrix $\mathbf{\Theta}$. The identification of the block structure, meaning the estimation of the number of clusters $K$ and cluster composition, is done as follows. Every variable is initialized in its own cluster (i.e. $\hat{K}=p$). For a fixed value of the aggregation penalty $\lambda_{c}$ and initial estimate at $\widehat{\mathbf{\Theta}}$, the clusters that are eligible for fusion are determined by computing the distances $d_{jj^{\prime}}(\widehat{\mathbf{\Theta}})$ for every variable pair $j,j^{\prime}$. If this distance reduces to zero, the two corresponding variables $j$ and $j^{\prime}$ are eligible for fusion. The number of clusters $\hat{K}$ and cluster composition are then updated accordingly. Given this updated block structure, objective function (5) is optimized for every block/cluster $k=1,\ldots,\hat{K}$.
@@ 2.3.1 Outline of the Cyclic Block Coordinate Descent Algorithm
Cluster fusions. The first step is to verify whether cluster $k$ is close enough to another cluster to warrant their fusion for estimate $\widehat{\mathbf{\Theta}}$ (in a given iteration of the algorithm). Let $\mathcal{C}_{k}$ denote the set of variables belonging to cluster $k$, analogously $\mathcal{C}_{\ell}$ for cluster $\ell$. Then, cluster $k$ is fused with cluster $\ell$ if $d_{\mathcal{C}_{k}\mathcal{C}_{\ell}}(\widehat{\mathbf{\Theta}})$, which measures the distance between blocks $\widehat{\mathbf{\Theta}}_{k}$ and $\widehat{\mathbf{\Theta}}_{\ell}$, is smaller than some user-defined threshold $\varepsilon_{f}$ representing the minimum required similarity for fusion. In case of multiple candidates for which this applies, cluster $\ell$ is chosen as $\ell=\operatorname*{argmin}_{\ell^{\prime}}d_{\mathcal{C}_{k}\mathcal{C}_{\ell^{\prime}}}(\widehat{\mathbf{\Theta}})$. If a fusion is performed, it advances to the next cluster. If no eligible fusion candidate is found in the first step, we proceed to the second step where we update the parameter estimates for cluster $k$.
@@
Finally, we determine the optimal step size $s^{*}$ given the descent direction using an inexact line search. The reason for augmenting the Newton descent direction with a line search is twofold. First, the minimization of $L(\mathbf{\Theta}_{k})$ is constrained by the restriction that $\mathbf{\Theta}$ should be positive definite. Hence, a step size should be chosen that ensures the update to adhere this restriction. In Appendix B.1, we derive the conditions under which the optimal step size can be computed that adheres to the positive definiteness restriction. A second reason to opt for a line search is the fact that the Hessian may not be well-behaved due to the presence of the $\ell_{2}$-norm in the clusterpath penalty. If, for the current value $\widehat{\mathbf{\Theta}}$, at least one of the distances $d_{\mathcal{C}_{k}\mathcal{C}_{\ell}}(\widehat{\mathbf{\Theta}})$ is close to zero, $L(\widehat{\mathbf{\Theta}}_{k})$ is not locally smooth. As a result, the Hessian is not well-behaved and the favorable convergence properties of Newton’s method do not hold. A line search for the optimal step size guarantees that the update does not increase the objective function.
@@ 2.3.2 Weight Matrix
An important element of the clusterpath estimator is the weight matrix $\mathbf{W}$ that contains information about the preferences of clustering variables $j$ and $j^{\prime}$ through the weight $w_{jj^{\prime}}$. A large value for $w_{jj^{\prime}}$ incentivizes clustering of the corresponding variables, but it does not guarantee this. In the convex clustering literature, weights are typically based on the squared distances between data points, with performance improving when using $k$-nearest neighbors to assign zero weights to nonneighbor pairs (e.g., Chen et al.,, 2015; Chi & Lange,, 2015). Following these standard practices, we scale the squared distances by the mean squared distance and take
@@
A potential issue with relying solely on $k$-nearest neighbors for the construction of $\mathcal{E}$ is the occurrence of disconnected subgroups. To alleviate this issue, we add variable pairs to $\mathcal{E}$ using the procedure described in Touw et al., (2022) until all disconnected subgroups are eliminated. This approach effectively integrates the $k$-nearest neighbors structure with a minimum spanning tree to obtain a weight matrix corresponding to a connected graph.
@@ 2.3.4 Refitted CGGM
While minimization of the objective function in (2) yields a possibly sparse and/or clustered $\widehat{\mathbf{\Theta}}$, the CGGM estimate may be biased due to the shrinkage of the distances between different clusters and the shrinkage of the entries in the precision matrix to zero. Following Wilms & Bien, (2022), we therefore also consider a refitted version of CGGM. First, CGGM is used to obtain a clustering and sparsity pattern in $\widehat{\mathbf{\Theta}}$. We then re-estimate the precision matrix $\mathbf{\Theta}$ constrained by the obtained clustering and sparsity by maximizing the likelihood subject to the clustering and sparsity constraint. By re-estimating the precision matrix, we aim to capitalize on the clustering- and sparsity-inducing capabilities of CGGM while refining parameter estimates for improved accuracy.
@@ 3.1 Simulation Designs
For most of the simulations, we follow the designs of Wilms & Bien, (2022) together with additional variations. In all settings, we sample data from a multivariate normal distribution with mean zero and covariance $\mathbf{\Sigma}=\mathbf{\Theta}^{-1}$, and we repeat the process 100 times.
@@
Increasing the number of variables. We focus on the chain design as in Wilms & Bien, (2022) and vary the number of variables $p\in\{15,30,60,120\}$. We set the corresponding number of observations $n\in\{120,240,480,960\}$ such that the ratio $n/p$ is constant. The number of clusters is kept fixed at $K=3$.
Increasing the number of clusters. We again focus on the chain design and vary the number of clusters $K\in\{3,5,6,10\}$ while keeping the number of observations and variables fixed at $n=240$ and $p=30$, respectively, as in Wilms & Bien, (2022).
@@
Blockdiagonal structure. In these designs, the precision matrix $\mathbf{\Theta}$ does not exhibit a variable clustering structure, but a blockdiagonal structure inspired by the simulation design of Pircalabelu & Claeskens, (2020). We use the same dimensions as in the baseline design and $K=3$ blocks on the diagonal. Outside of the diagonal blocks, edges between variables are drawn with probability $\pi=0.1$. The diagonal values of $\mathbf{\Theta}$ are set to 1, the off-diagonal elements in the diagonal blocks are set to 0.5, and non-zero elements outside the diagonal blocks are set to 0.25. If such a randomly drawn precision matrix is not positive semi-definite, we repeat this process until a positive semi-definite precision matrix is obtained. We consider both a balanced design with $p_{1}=p_{2}=p_{3}=5$, and an unbalanced design with $p_{1}=3$, $p_{2}=5$, and $p_{3}=7$ (see right two panels in bottom row of Figure 2).
@@ 3.2 Methods and Evaluation Criteria
We apply CGGM with and without the parameter re-estimation step described in Section 2.3.4, which we refer to as CGGM-raw and CGGM-refit. We thereby set the convergence tolerance $\varepsilon_{c}=10^{-7}$ and the maximum number of iterations to $t_{\text{max}}=5000$. For comparison, we include the tree-aggregated graphical lasso (TAGL) of Wilms & Bien, (2022), the community-based group graphical lasso (ComGGL) of Pircalabelu & Claeskens, (2020), the graphical lasso (GL) (Friedman et al.,, 2008), and the inverse of the sample covariance matrix ($\mathbf{S}^{-1}$).
TAGL performs node aggregation based on side-information in the form of a tree-based variable hierarchy, which needs to be specified a priori. We generate an ideal and a realistic tree hierarchy as described in Wilms & Bien, (2022) (see Figure 5 in their paper for an illustration). As both trees contain the true variable clustering, we also generate a misspecified tree in the same manner as the realistic tree, except that with probability 0.1, each variable is incorrectly assigned to the subsequent cluster (or the first cluster if the variable in fact belongs to the last cluster).
@@
Concerning tuning parameter selection, we apply 3-fold cross-validation. For the weight matrix in CGGM, we use candidate values $k\in\{1,3,5\}$ for the number of neighbors and we set $\phi=1$ to keep the computational burden low. We verified that for this choice of $\phi$, the weights showed sufficient variation over the interval $(0,1]$ to stimulate variable clustering. Moreover, we select candidate value for the regularization parameters $\lambda_{c}$ and $\lambda_{s}$ as described in Section 2.3.3. For TAGL, we first perform a binary search for the smallest value of the aggregation parameter that aggregates all variables into one cluster (for each of the tree hierarchies) while keeping the sparsity parameter at 0, and we determine the smallest value of the sparsity parameter that removes all edges while keeping the aggregation parameter at 0 (corresponding to the graphical lasso). Ten candidate values for the aggregation and sparsity parameters are then obtained as fractions of those aforementioned values, where the fractions vary from 0 to 1, with each step being twice as large as the previous step. For ComGGL, we similarly set ten candidate values for the grouping and sparsity parameters as fractions of maximum values. While we fix the maximum value of the grouping parameter at 1, we again set the maximum value of the sparsity parameter as the smallest value that removes all edges in the graphical lasso. As in Pircalabelu & Claeskens, (2020), we set the balancing parameter to 1. For the graphical lasso, we choose ten candidate values for the sparsity parameter in the same manner.
We evaluate the methods in terms of their estimation accuracy, aggregation, and sparsity recognition performance. Regarding estimation accuracy, we compute the Frobenius norm $\|\mathbf{\Theta}-\mathbf{\hat{\Theta}}\|_{F}$. For aggregation performance, we report the estimated number of clusters $\hat{K}$ as well as the adjusted Rand index (ARI) (Hubert & Arabie,, 1985) of the obtained clustering of variables. Note that $\hat{K}$ directly follows from the obtained clustering for the optimal values of the tuning parameters from cross-validation. For sparsity recognition, we report the false positive and false negative rates (FPR and FNR, respectively), where the FPR reports the fraction of true zero elements of the precision matrix that are estimated as nonzero and the FNR gives the fraction of true nonzero elements of the precision matrix that are estimated as zero.
@@ 3.3 Results
Baseline simulation designs. In terms of estimation accuracy and aggregation, CGGM-raw and CGGM-refit perform similar to TAGL-ideal, outperform TAGL-realistic in some designs, and outperform and TAGL-misspecified in all designs (see Figure 3). Regarding sparsity recognition, one variant of CGGM generally outperforms all variants of TAGL, depending on the setting. The refitting step of CGGM is advantageous in the three block designs but disadvantageous in the unstructured design. For the latter, the refitting step often aggregates all variables into one cluster, likely due to the high degree of sparsity in the unstructured design. This pooling of estimates for the large number of zero elements reduces noise, potentially improving the out-of-sample cross-validation score in Equation (7), despite introducing some bias in the relatively small number of nonzero elements. Conversely, CGGM-raw allows for shrinkage between elements without aggregation, thus improving estimation performance for designs in which many off-diagonal elements share the same value without belonging to the same cluster. While TAGL-misspecified often selects the correct number of clusters, the ARI reveals that it struggles to accurately classify the variables within those clusters due to the misspecified tree. This pattern persists across most of the simulation designs. Note that across all simulations, ComGGL detects a single community. This should, however, not be interpreted as a block-structured precision matrix estimate with a single cluster in the sense of equation (2.1), as ComGGL allows entries of the precision matrix estimate within the same community to vary.
@@ 4 Estimation of a Clustered Covariance Matrix
Existing paradigms for regularized estimation either induce an appropriate simplicity structure on the precision matrix or the covariance matrix, and the induced simplicity structure is, in general, not maintained when taking the inverse of the object of interest (e.g., Yao & Allen,, 2019, Pircalabelu & Claeskens,, 2020, and Wilms & Bien,, 2022). The proposed CGGM estimator connects both paradigms since it retains the found block structure when taking the inverse of the obtained estimate (Section 2.1). If model (2.1) holds for the precision matrix $\mathbf{\Theta}$, it may therefore be reasonable to obtain a clustered estimate of the covariance matrix $\mathbf{\Sigma}$ by taking $\mathbf{\hat{\Sigma}}=\mathbf{\hat{\Theta}}^{-1}$, where $\mathbf{\hat{\Theta}}$ denotes the CGGM estimate of the precision matrix. While the induced shared block structure in the precision and covariance matrix is a distinct feature of our proposal, adherence to an exact block structure as in model (2.1) may, however, be a strong assumption in practice. If this model only holds approximately with somewhat similar values within the blocks of $\mathbf{\Theta}$, we have seen in the simulations that estimating a clustered precision matrix with CGGM is nevertheless beneficial. But if an approximate block structure is present in the covariance matrix $\mathbf{\Sigma}$, the structure in $\mathbf{\Theta}$ is much more noisy with larger variability between the elements in the corresponding blocks, see Figure 5. In such settings, first estimating a clustered precision matrix with CGGM and then taking the inverse may not succeed in finding the correct structure.
@@
$-\log|\mathbf{\Sigma}^{-1}|+\operatorname*{tr}\mathbf{S}\mathbf{\Sigma}^{-1}+\lambda\mathcal{P}(\mathbf{\Sigma}).$

### v3

@@ 2.3 Cyclic Block Coordinate Descent Algorithm
We develop a cyclic block coordinate descent algorithm, tailored to minimizing objective function (2). [^1] One could opt for alternative algorithms such as the alternating directions method of multipliers (ADMM), provided it is tailored towards solving problem (2); as for instance \@BBOPcite\@BAP\@BBNWilms and Bien (2022)\@BBCP do for their method. However, while all the subproblems of the ADMM by \@BBOPcite\@BAP\@BBNWilms and Bien (2022)\@BBCP are solvable in closed-form, this would not be the case for the subproblem corresponding to our proposed aggregation penalty. This reason, in combination with the natural link between the block structure in $\bm{\Theta}$ and the block structure of the cyclic block coordinate descent algorithm further explains our choice for the latter. We opt for cyclic block coordinate descent since the blocks are naturally formed by the clusters $k=1,\ldots,K$. By exploiting this block structure in $\mathbf{\Theta}$, our algorithm builds on efficient expressions of the objective function, which allows us to minimize optimization problem (2) in a computationally efficient way.
@@
[…]
@@ 2.3.1 Outline of the Cyclic Block Coordinate Descent Algorithm
[…]
@@
[…]
@@ 2.3.2 Weight Matrix
An important element of the clusterpath estimator is the weight matrix $\mathbf{W}$ that contains information about the preferences of clustering variables $j$ and $j^{\prime}$ through the weight $w_{jj^{\prime}}$. A large value for $w_{jj^{\prime}}$ incentivizes clustering of the corresponding variables, but it does not guarantee this. In the convex clustering literature, weights are typically based on the squared distances between data points, with performance improving when using $k$-nearest neighbors to assign zero weights to nonneighbor pairs \@BBOPcitep\@BAP\@BBN(e.g., Chen et al., 2015; Chi and Lange, 2015)\@BBCP. Following these standard practices, we scale the squared distances by the mean squared distance and take
@@
A potential issue with relying solely on $k$-nearest neighbors for the construction of $\mathcal{E}$ is the occurrence of disconnected subgroups. To alleviate this issue, we add variable pairs to $\mathcal{E}$ using the procedure described in \@BBOPcite\@BAP\@BBNTouw et al. (2022)\@BBCP until all disconnected subgroups are eliminated. This approach effectively integrates the $k$-nearest neighbors structure with a minimum spanning tree to obtain a weight matrix corresponding to a connected graph.
@@ 2.3.4 Refitted CGGM
While minimization of the objective function in (2) yields a possibly sparse and/or clustered $\widehat{\mathbf{\Theta}}$, the CGGM estimate may be biased due to the shrinkage of the distances between different clusters and the shrinkage of the entries in the precision matrix to zero. Following \@BBOPcite\@BAP\@BBNWilms and Bien (2022)\@BBCP, we therefore also consider a refitted version of CGGM. First, CGGM is used to obtain a clustering and sparsity pattern in $\widehat{\mathbf{\Theta}}$. We then re-estimate the precision matrix $\mathbf{\Theta}$ constrained by the obtained clustering and sparsity by maximizing the likelihood subject to the clustering and sparsity constraint. By re-estimating the precision matrix, we aim to capitalize on the clustering- and sparsity-inducing capabilities of CGGM while refining parameter estimates for improved accuracy.
@@ 3.1 Simulation Designs
For most of the simulations, we follow the designs of \@BBOPcitet\@BAP\@BBNWilms and Bien (2022)\@BBCP together with additional variations. In all settings, we sample data from a multivariate normal distribution with mean zero and covariance $\mathbf{\Sigma}=\mathbf{\Theta}^{-1}$, and we repeat the process 100 times.
@@
Increasing the number of variables. We focus on the chain design as in \@BBOPcitet\@BAP\@BBNWilms and Bien (2022)\@BBCP and vary the number of variables $p\in\{15,30,60,120\}$. We set the corresponding number of observations $n\in\{120,240,480,960\}$ such that the ratio $n/p$ is constant. The number of clusters is kept fixed at $K=3$.
Increasing the number of clusters. We again focus on the chain design and vary the number of clusters $K\in\{3,5,6,10\}$ while keeping the number of observations and variables fixed at $n=240$ and $p=30$, respectively, as in \@BBOPcitet\@BAP\@BBNWilms and Bien (2022)\@BBCP.
@@
Blockdiagonal structure. In these designs, the precision matrix $\mathbf{\Theta}$ does not exhibit a variable clustering structure, but a blockdiagonal structure inspired by the simulation design of \@BBOPcitet\@BAP\@BBNPircalabelu and Claeskens (2020)\@BBCP. We use the same dimensions as in the baseline design and $K=3$ blocks on the diagonal. Outside of the diagonal blocks, edges between variables are drawn with probability $\pi=0.1$. The diagonal values of $\mathbf{\Theta}$ are set to 1, the off-diagonal elements in the diagonal blocks are set to 0.5, and non-zero elements outside the diagonal blocks are set to 0.25. If such a randomly drawn precision matrix is not positive semi-definite, we repeat this process until a positive semi-definite precision matrix is obtained. We consider both a balanced design with $p_{1}=p_{2}=p_{3}=5$, and an unbalanced design with $p_{1}=3$, $p_{2}=5$, and $p_{3}=7$ (see right two panels in bottom row of Figure 2).
@@ 3.2 Methods and Evaluation Criteria
We apply CGGM with and without the parameter re-estimation step described in Section 2.3.4, which we refer to as CGGM-raw and CGGM-refit. We thereby set the convergence tolerance $\varepsilon_{c}=10^{-7}$ and the maximum number of iterations to $t_{\text{max}}=5000$. For comparison, we include the tree-aggregated graphical lasso (TAGL) of \@BBOPcitet\@BAP\@BBNWilms and Bien (2022)\@BBCP, the community-based group graphical lasso (ComGGL) of \@BBOPcitet\@BAP\@BBNPircalabelu and Claeskens (2020)\@BBCP, the graphical lasso (GL) \@BBOPcitep\@BAP\@BBN(Friedman et al., 2008)\@BBCP, and the inverse of the sample covariance matrix ($\mathbf{S}^{-1}$).
TAGL performs node aggregation based on side-information in the form of a tree-based variable hierarchy, which needs to be specified a priori. We generate an ideal and a realistic tree hierarchy as described in \@BBOPcitet\@BAP\@BBNWilms and Bien (2022)\@BBCP (see Figure 5 in their paper for an illustration). As both trees contain the true variable clustering, we also generate a misspecified tree in the same manner as the realistic tree, except that with probability 0.1, each variable is incorrectly assigned to the subsequent cluster (or the first cluster if the variable in fact belongs to the last cluster).
@@
Concerning tuning parameter selection, we apply 3-fold cross-validation. For the weight matrix in CGGM, we use candidate values $k\in\{1,3,5\}$ for the number of neighbors and we set $\phi=1$ to keep the computational burden low. We verified that for this choice of $\phi$, the weights showed sufficient variation over the interval $(0,1]$ to stimulate variable clustering. Moreover, we select candidate value for the regularization parameters $\lambda_{c}$ and $\lambda_{s}$ as described in Section 2.3.3. For TAGL, we first perform a binary search for the smallest value of the aggregation parameter that aggregates all variables into one cluster (for each of the tree hierarchies) while keeping the sparsity parameter at 0, and we determine the smallest value of the sparsity parameter that removes all edges while keeping the aggregation parameter at 0 (corresponding to the graphical lasso). Ten candidate values for the aggregation and sparsity parameters are then obtained as fractions of those aforementioned values, where the fractions vary from 0 to 1, with each step being twice as large as the previous step. For ComGGL, we similarly set ten candidate values for the grouping and sparsity parameters as fractions of maximum values. While we fix the maximum value of the grouping parameter at 1, we again set the maximum value of the sparsity parameter as the smallest value that removes all edges in the graphical lasso. As in \@BBOPcitet\@BAP\@BBNPircalabelu and Claeskens (2020)\@BBCP, we set the balancing parameter to 1. For the graphical lasso, we choose ten candidate values for the sparsity parameter in the same manner.
We evaluate the methods in terms of their estimation accuracy, aggregation, and sparsity recognition performance. Regarding estimation accuracy, we compute the Frobenius norm $\|\mathbf{\Theta}-\mathbf{\hat{\Theta}}\|_{F}$. For aggregation performance, we report the estimated number of clusters $\hat{K}$ as well as the adjusted Rand index (ARI) \@BBOPcitep\@BAP\@BBN(Hubert and Arabie, 1985)\@BBCP of the obtained clustering of variables. Note that $\hat{K}$ directly follows from the obtained clustering for the optimal values of the tuning parameters from cross-validation. For sparsity recognition, we report the false positive and false negative rates (FPR and FNR, respectively), where the FPR reports the fraction of true zero elements of the precision matrix that are estimated as nonzero and the FNR gives the fraction of true nonzero elements of the precision matrix that are estimated as zero.
@@ 3.3 Results
Baseline simulation designs. In terms of estimation accuracy and aggregation, CGGM-raw and CGGM-refit perform similar to TAGL-ideal, outperform TAGL-realistic in some designs, and outperform and TAGL-misspecified in all designs (see Figure 3). Regarding sparsity recognition, one variant of CGGM generally outperforms all variants of TAGL, depending on the setting. The refitting step of CGGM is advantageous in the three block designs but disadvantageous in the unstructured design. For the latter, the refitting step often yields only one or two clusters, likely due to the high degree of sparsity in the unstructured design. This pooling of estimates for the large number of zero elements reduces noise, potentially improving the out-of-sample cross-validation score in Equation (7), despite introducing some bias in the relatively small number of nonzero elements. Conversely, CGGM-raw allows for shrinkage between elements without aggregation, thus improving estimation performance for designs in which many off-diagonal elements share the same value without belonging to the same cluster. While TAGL-misspecified often selects the correct number of clusters, the ARI reveals that it struggles to accurately classify the variables within those clusters due to the misspecified tree. This pattern persists across most of the simulation designs. Note that across all simulations, ComGGL detects a single community. This should, however, not be interpreted as a block-structured precision matrix estimate with a single cluster in the sense of Equation (2.1), as ComGGL allows entries of the precision matrix estimate within the same community to vary.
@@ 4 Estimation of a Clustered Covariance Matrix
Existing paradigms for regularized estimation either induce an appropriate simplicity structure on the precision matrix or the covariance matrix, and the induced simplicity structure is, in general, not maintained when taking the inverse of the object of interest (e.g., \@BBOPcitep\@BAP\@BBNYao and Allen, 2019\@BBCP, \@BBOPcitep\@BAP\@BBNPircalabelu and Claeskens, 2020\@BBCP, and \@BBOPcitep\@BAP\@BBNWilms and Bien, 2022\@BBCP). The proposed CGGM estimator connects both paradigms since it retains the found block structure when taking the inverse of the obtained estimate (Section 2.1). If model (2.1) holds for the precision matrix $\mathbf{\Theta}$, it may therefore be reasonable to obtain a clustered estimate of the covariance matrix $\mathbf{\Sigma}$ by taking $\mathbf{\hat{\Sigma}}=\mathbf{\hat{\Theta}}^{-1}$, where $\mathbf{\hat{\Theta}}$ denotes the CGGM estimate of the precision matrix. While the induced shared block structure in the precision and covariance matrix is a distinct feature of our proposal, adherence to an exact block structure as in model (2.1) may, however, be a strong assumption in practice. If this model only holds approximately with somewhat similar values within the blocks of $\mathbf{\Theta}$, we have seen in the simulations that estimating a clustered precision matrix with CGGM is nevertheless beneficial. But if an approximate block structure is present in the covariance matrix $\mathbf{\Sigma}$, the structure in $\mathbf{\Theta}$ is much more noisy with larger variability between the elements in the corresponding blocks, see Figure 5. In such settings, first estimating a clustered precision matrix with CGGM and then taking the inverse may not succeed in finding the correct structure.
@@
$\log|\mathbf{\Sigma}^{-1}|+\operatorname*{tr}\mathbf{S}\mathbf{\Sigma}^{-1}+\lambda\mathcal{P}(\mathbf{\Sigma}).$